# Understanding Learning Dynamics Through Structured Representations

### v1

@@ II-A Preliminaries and Setup
$x^{(l)}=S^{(l)}W^{(l)}x^{(l-1)}+C^{(l)}(x^{(l-1)}),$ (1)
where $S^{(l)}$ is a shaping operator that constrains or modulates $W^{(l)}$, and $C^{(l)}$ is a learnable correction function. This decomposition generalizes standard affine layers by allowing explicit architectural constraints to act before optimization begins.
@@ II-B Gradient Flow and Spectral Conditioning
II-C Directional Selectivity and Functional Smoothing
Structured pathways can act as low-pass filters in functional space, suppressing rapid changes and focusing learning capacity on smoother or more coherent directions. This effect is especially useful in overparameterized models, where unconstrained directions may absorb noise or reinforce spurious correlations.
We define a directional selectivity index (DSI) based on the projection of learned updates onto the dominant eigenspaces of $S^{(l)}$:
$\text{DSI}^{(l)}=\frac{\|\mathbb{P}_{\text{dom}}\nabla_{W^{(l)}}\mathcal{L}\|}{\|\nabla_{W^{(l)}}\mathcal{L}\|},$ (3)
[…]
II-D Corrective Paths as Dynamic Modulators
The correction term $C^{(l)}$ plays a dual role: it preserves model expressivity and absorbs residual error that structured components cannot capture. Unlike architectural mechanisms such as batch normalization or residual connections, $C^{(l)}$ does not merely pass identity or rescale activations—it actively learns to modulate the shaped transformation.
This dynamic modulation can be interpreted as a kind of learned response function that adapts to compensate for rigidity in $S^{(l)}W^{(l)}$. As such, the interaction between structured and corrective paths forms a two-way system: one enforces stable propagation, the other flexibly recovers expressivity.
II-E Implications for Depth and Robustness
One of the primary benefits of this formulation is improved depth scalability. By ensuring that each transformation respects spectral constraints and directional smoothness, the model can be stacked deeper without requiring skip connections or aggressive normalization. Moreover, the coupling with $C^{(l)}$ provides a built-in mechanism for local adaptation, mitigating the risks of structural bias or underfitting.
@@ IV-A Structural Bias as Inductive Prior
IV-B Expressive Compensation via Corrective Path
While $S^{(l)}W^{(l)}$ enforces constraints, the correction path $C^{(l)}$ acts as a compensatory mechanism that restores expressivity. This dual-channel architecture allows the network to remain expressive without relying entirely on unconstrained parameters. Empirical results (see Section VI) show that ablating $C^{(l)}$ leads to underfitting, while removing $S^{(l)}$ increases variance and sensitivity to noise.
This controlled division of labor creates a natural expressive tradeoff: the structured path governs global behavior, while the correction path captures local deviation. Unlike hand-crafted bottlenecks or attention masks, this tradeoff emerges from the design itself.
IV-C Generalization Bounds and Hypothesis Class
@@
IV-D Emergent Modularity and Compositional Behavior
@@
IV-E A Generalization Bound via Structured Composition
@@
=$\mathcal{H}=\{h(x)=SWx+\phi(x)\mid W\in\mathbb{R}^{d\times d},\phi\in\mathcal{H}_{C}\}.$ (5)
@@
Let $\mathcal{H}$ be the structured architecture class as above, and suppose the loss function $\ell(h(x),y)$ is $L$-Lipschitz and bounded in [0,1][0,1][ 0 , 1 ]. Then, with probability at least $1-\delta$ over a sample of size $n$, every $h\in\mathcal{H}$ satisfies:
@@
We begin with a generalization bound for Lipschitz losses (see Bartlett and Mendelson, 2002). Since $\ell$ is $L$-Lipschitz and bounded in [0,1][0,1][ 0 , 1 ], with probability at least $1-\delta$, for all $h\in\mathcal{H}$:
@@
$=\frac{1}{n}\mathbb{E}_{\sigma}\left[\sup_{\|W\|_{F}\leq B}\left\langle W,\sum_{i=1}^{n}\sigma_{i}S^{\top}(ex_{i}^{\top})\right\rangle_{F}\right]=\frac{B}{n}\mathbb{E}_{\sigma}\left[\left\|\sum_{i=1}^{n}\sigma_{i}S^{\top}(ex_{i}^{\top})\right\|_{F}\right].$
@@ V Convergence of Structured Recursive Dynamics
Theorem. Assume $SW$ and $\phi$ are Lipschitz continuous with constants $L_{1}$, $L_{2}$ such that:
@@
Proof. Let $F(x)=SWx+\phi(x)$. Using the triangle inequality:
@@ V-A Examples and Interpretations
Theorem V guarantees convergence under a simple condition on the combined Lipschitz constants of the structured linear path and the correction component. We now illustrate concrete instances where this condition is satisfied.
@@ V-A1 Example 1: Orthogonal Structure with Bounded Nonlinearity
If $\phi(x)=\alpha\cdot\tanh(Bx)$ with $\|\alpha B\|\leq 0.2$, then $L_{2}\leq 0.2$ and total Lipschitz constant becomes $L_{1}+L_{2}=1.2$. This exceeds 1, so the contraction condition is not satisfied. However, with a small scaling of $W$ (e.g., $\|W\|=0.7$), the combined Lipschitz constant becomes 0.90.90.9, ensuring convergence.
@@ V-A2 Example 2: Low-Rank Projection and Shallow Residual
Suppose $S$ is a fixed low-rank projection (e.g., a Laplacian or spectral filter), with induced norm $\|S\|=0.5$, and $W$ is arbitrary. Then $\|SW\|\leq 0.5\|W\|$. If $W$ is initialized with norm $\|W\|\leq 1.0$, we get $L_{1}\leq 0.5$.
@@ V-A3 Example 3: Diagonal Scaling with Smooth Activation
Let $\phi(x)=\sigma(Bx)$ with $\sigma$ being a softsign or ELU activation. Suppose $\|B\|=0.5$ and the effective Lipschitz constant of $\phi$ is bounded by 0.40.40.4. Then again $L_{1}+L_{2}=0.8$, and convergence follows.
@@
VI Empirical Support for Structural Hypotheses
This section provides empirical illustrations that support the theoretical constructs introduced above. Rather than focusing on benchmark performance or exhaustive comparisons, we aim to validate how structured transformations influence learning behavior, convergence, and representational consistency under varying conditions. All experiments are deliberately chosen to isolate specific aspects of the proposed theoretical framework.
VI-A Sensitivity to Input Alignment
@@
Observation. In the aligned setting, structured networks converge significantly faster and exhibit lower gradient variance. In the randomized setting, convergence is slower for both, but the structured network still stabilizes earlier.
@@
VI-B Projection Incompleteness and Correction Load
@@
VI-C Stability Under Gradient Perturbation
@@
VI-D Pathway Decoupling Analysis
@@
VII Task-Level Comparison on Fashion-MNIST
To examine how the proposed structured-corrective architecture behaves in practice, we compare it against a standard multilayer perceptron (MLP) on a real-world image classification task. This section does not aim to achieve state-of-the-art accuracy, but instead focuses on emergent training behavior and reliability across runs.
VII-A Setup
We consider the Fashion-MNIST dataset with two-layer networks using either (i) standard MLP layers or (ii) PGNN blocks as introduced earlier. Both models are trained for 20 epochs using the Adam optimizer and identical hyperparameters. The PGNN model employs an identity-based shaping operator and a learned correction path, while the MLP baseline is purely learned.
VII-B Accuracy Comparison
Figure 1 shows both training and test accuracy over epochs. PGNN achieves comparable final performance to MLP but exhibits smoother learning dynamics and slightly faster convergence in earlier epochs.
Figure 1: Training and test accuracy comparison between MLP and PGNN on Fashion-MNIST.
VII-C Loss Dynamics
Figure 2 shows the cross-entropy loss curves for both models. Notably, the PGNN model consistently maintains a tighter gap between train and test loss, suggesting better generalization behavior.
Figure 2: Cross-entropy loss during training for MLP and PGNN models.
VII-D Initialization Sensitivity
To probe robustness, we run each model over 5 random seeds and compute mean and standard deviation of test accuracy at each epoch. Results in Figure 3 show that PGNN exhibits more stable convergence and reduced variance over seeds, particularly in the early and middle epochs.
Figure 3: Test accuracy (mean ± std) across 5 random seeds. PGNN shows reduced initialization sensitivity.
VII-E Interpretation
@@
VIII Conclusion

### v2

@@ II-A Preliminaries and Setup
$x^{(l)}=\underbrace{\mathcal{S}^{(l)}W^{(l)}x^{(l-1)}}_{\text{Structure-Aware Path}}+\underbrace{\mathcal{C}^{(l)}(x^{(l-1)})}_{\text{Adaptive Correction Path}},$ (1)
where $\mathcal{S}^{(l)}$ is a shaping operator that constrains or modulates $W^{(l)}$, and $\mathcal{C}^{(l)}$ is a learnable correction function. This decomposition generalizes standard affine layers by allowing explicit architectural constraints to act before optimization begins. We refer to this structured–corrective architecture as a Physics-Guided Neural Network (PGNN).
@@ II-B Gradient Flow and Spectral Conditioning
Directional bias
Structured paths bias gradient energy toward dominant modes of $\mathcal{S}^{(l)}$. We evidence this via Jacobian-spectrum and gradient-norm diagnostics (Sec. VIII) and model-derived CKA/Subspace-Overlap (Sec. VI).
II-C Corrective Paths as Dynamic Modulators
The correction term $\mathcal{C}^{(l)}$ plays a dual role: it preserves model expressivity and absorbs residual error that structured components cannot capture. Unlike architectural mechanisms such as batch normalization or residual connections, $\mathcal{C}^{(l)}$ does not merely pass identity or rescale activations—it actively learns to modulate the shaped transformation.
This dynamic modulation can be interpreted as a kind of learned response function that adapts to compensate for rigidity in $\mathcal{S}^{(l)}W^{(l)}$. As such, the interaction between structured and corrective paths forms a two-way system: one enforces stable propagation, the other flexibly recovers expressivity.
II-D Implications for Depth and Robustness
One of the primary benefits of this formulation is improved depth scalability. By ensuring that each transformation respects spectral constraints and directional smoothness, the model can be stacked deeper without requiring skip connections or aggressive normalization. Moreover, the coupling with $\mathcal{C}^{(l)}$ provides a built-in mechanism for local adaptation, mitigating the risks of structural bias or underfitting.
@@ IV-A Structural Bias as Inductive Prior
IV-B Generalization Bounds and Hypothesis Class
@@
IV-C Emergent Modularity and Compositional Behavior
@@
IV-D A Generalization Bound via Structured Composition
@@
$\mathcal{H}=\{h(x)=SWx+\phi(x)\mid W\in\mathbb{R}^{d\times d},\phi\in\mathcal{H}_{C}\}.$ (4)
@@
Let $\mathcal{H}$ be the structured architecture class as above, and suppose the loss function $\ell(h(x),y)$ is $L$-Lipschitz and bounded in $[0,1]$. Then, with probability at least $1-\delta$ over a sample of size $n$, every $h\in\mathcal{H}$ satisfies:
@@
We begin with a generalization bound for Lipschitz losses (see Bartlett and Mendelson, 2002). Since $\ell$ is $L$-Lipschitz and bounded in $[0,1]$, with probability at least $1-\delta$, for all $h\in\mathcal{H}$:
@@
$\frac{1}{n}\mathbb{E}_{\sigma}\left[\sup_{\|W\|_{F}\leq B}\left\langle W,\sum_{i=1}^{n}\sigma_{i}S^{\top}(ex_{i}^{\top})\right\rangle_{F}\right]=\frac{B}{n}\mathbb{E}_{\sigma}\left[\left\|\sum_{i=1}^{n}\sigma_{i}S^{\top}(ex_{i}^{\top})\right\|_{F}\right].$
@@ V Convergence of Structured Recursive Dynamics
Theorem.
Theorem 2.
Assume $SW$ and $\phi$ are Lipschitz continuous with constants $L_{1}$, $L_{2}$ such that:
@@
Proof.
Let $F(x)=SWx+\phi(x)$. Using the triangle inequality:
@@ V-A Examples and Interpretations
Theorem 2 guarantees convergence under a simple condition on the combined Lipschitz constants of the structured linear path and the correction component. We now illustrate concrete instances where this condition is satisfied.
@@ V-A1 Example 1: Orthogonal Structure with Bounded Nonlinearity
If $\phi(x)=\alpha\cdot\tanh(Bx)$ with $\|\alpha B\|\leq 0.2$, then $L_{2}\leq 0.2$ and total Lipschitz constant becomes $L_{1}+L_{2}=1.2$. This exceeds 1, so the contraction condition is not satisfied. However, with a small scaling of $W$ (e.g., $\|W\|=0.7$), the combined Lipschitz constant becomes $0.9$, ensuring convergence.
@@ V-A2 Example 2: Low-Rank Projection and Shallow Residual
Suppose $S$ is a fixed low-rank projection, e.g., a Laplacian or spectral filter, with induced norm $\|S\|=0.5$, and $W$ is arbitrary. Then $\|SW\|\leq 0.5\|W\|$. If $W$ is initialized with norm $\|W\|\leq 1.0$, we get $L_{1}\leq 0.5$.
@@ V-A3 Example 3: Diagonal Scaling with Smooth Activation
Let $\phi(x)=\sigma(Bx)$ with $\sigma$ being a softsign or ELU activation. Suppose $\|B\|=0.5$ and the effective Lipschitz constant of $\phi$ is bounded by $0.4$. Then again $L_{1}+L_{2}=0.8$, and convergence follows.
@@
VI Mechanistic Explanation of PGNN Dynamics
The architectural formulation and stability properties of PGNN have been introduced in prior work [1]. In this section, we focus on complementary diagnostics that reveal how PGNN’s structured pathways influence internal learning dynamics, generalization behavior, and convergence.
VI-A Layer-Wise Representation Consistency
To assess internal consistency, we compute the generalization gap at intermediate layers of PGNN and a standard MLP. For each layer, we measure the mean squared error (MSE) between training and validation activations. This analysis reveals that PGNN exhibits smaller gaps across layers, suggesting that its structured transformations promote more stable and generalizable internal representations. These results are visualized in Figure 1.
Figure 1: Layer-wise generalization gaps (MSE) for PGNN and MLP. PGNN shows reduced discrepancy between training and validation activations.
VI-B Subspace Geometry and Representational Overlap
We quantify representational structure with two standard, self-contained metrics.
Centered linear CKA [30]. Given a held-out batch, let $H_{\ell}\!\in\!\mathbb{R}^{n\times d}$ be centered activations at layer $\ell$ and $Z\!\in\!\mathbb{R}^{n\times m}$ be centered targets (or features). With $C=I-\tfrac{1}{n}\mathbf{1}\mathbf{1}^{\top}$,
$\mathrm{CKA}(H_{\ell},Z)=\frac{\langle CH_{\ell}H_{\ell}^{\top}C,\;CZZ^{\top}C\rangle_{F}}{\|CH_{\ell}H_{\ell}^{\top}C\|_{F}\;\|CZZ^{\top}C\|_{F}}.$
Subspace-Overlap (SOV). Let $U_{\text{tr}},U_{\text{ho}}\in\mathbb{R}^{d\times k}$ be orthonormal bases for the top-$k$ principal-component subspaces of the centered train and held-out activations at layer $\ell$. The overlap is the mean squared cosine of the principal angles:
$\mathrm{SOV}_{k}(H_{\ell}^{\text{tr}},H_{\ell}^{\text{ho}})=\tfrac{1}{k}\,\|U_{\text{tr}}^{\top}U_{\text{ho}}\|_{F}^{2}.$
We use k=16k{=}16 in all runs.
[…]
Figure 2: Model-derived CKA and Subspace-Overlap on a synthetic linear+noise classification task (d=16, 2 classes). Results over 5 seeds, single held-out batch; layers L1–L2.
L1 SOV $\uparrow$ L2 SOV $\uparrow$
MLP 0.934 $\pm$ 0.013 0.940 $\pm$ 0.012
PGNN 0.923 $\pm$ 0.018 0.951 $\pm$ 0.016
[…]
VI-C Equilibrium Interpretation
Unlike standard feedforward models, PGNN layers can be interpreted as recursive systems that converge toward stable representations. We simulate iterative application of PGNN blocks and measure convergence behavior using a contraction metric. These results are visualized in Figure 3.
Figure 3: Contraction metric over recursive PGNN iterations. The output stabilizes over time, indicating equilibrium behavior.
These diagnostics collectively demonstrate that PGNN’s structured design not only improves training stability but also induces smoother, more robust, and generalizable learning dynamics—complementing its architectural foundations without duplicating prior results.
Implementation details and experimental configurations for all diagnostics are provided in Section VII.
VI-D Synthetic Alignment Task (Formal Spec)
{specbox}
Goal. Quantify how closely hidden representations align with a known signal subspace using model-derived (real) activations.
Data. Input dim d=16d{=}16. Sample an orthonormal basis $U\!\in\!\mathbb{R}^{d\times d}$ (QR on a Gaussian matrix). For each input $x\!\sim\!\mathcal{N}(0,I_{d})$ define $z=U^{\top}x$. Targets are $y=Vz+\varepsilon$ with $V\!\in\!\mathbb{R}^{m\times d}$ (m=4m{=}4) and $\varepsilon\!\sim\!\mathcal{N}(0,\sigma^{2}I_{m})$, $\sigma{=}0.05$.
Models. Two hidden layers (width 64). Baseline MLP vs. PGNN with identical parameter budget:
$x^{(l)}=\mathcal{S}^{(l)}W^{(l)}x^{(l-1)}+\mathcal{C}^{(l)}\!\left(x^{(l-1)}\right),$
with $\mathcal{S}^{(l)}{=}I$ unless noted and $\mathcal{C}^{(l)}=\text{Linear}\!\rightarrow\!\text{ReLU}$ (two blocks total).
Training. Adam ($10^{-3}$), batch 128, $2000$ steps, $5$ seeds.
[…]
Model CKA $\uparrow$ Subspace-Overlap $\uparrow$ Transfer-Probe $R^{2}$ $\uparrow$
MLP 0.39 $\pm$ 0.02 0.11 $\pm$ 0.02 1.00 $\pm$ 0.00
PGNN 0.40 $\pm$ 0.02 0.13 $\pm$ 0.01 1.00 $\pm$ 0.00
Table II: Synthetic Alignment Task. Metrics computed from model-derived activations; 5 seeds, one held-out batch, best hidden layer per seed.
On this controlled task, PGNN shows a small but consistent gain in both CKA (+0.01) and SOV (+0.02) over the MLP, while the transfer probe saturates at $R^{2}{\approx}1.0$, as expected for the near-linear ground truth (see Table II).
VII Empirical Support for Structural Hypotheses
This section presents targeted diagnostics that isolate internal mechanisms of PGNN behavior—such as representation consistency, directional selectivity, and convergence dynamics. These analyses complement the broader empirical studies presented in Section VIII.
VII-A Experimental Setup
All experiments were implemented in PyTorch with fixed random seeds to ensure reproducibility. Synthetic datasets were generated using standard normal distributions, with input dimensions typically set to $d=16$ and batch sizes ranging from 100 to 200 depending on the task. For regression tasks, outputs were derived from low-dimensional signal components with optional Gaussian noise. For classification tasks, labels were computed from structured projections or latent mappings.
PGNN layers follow Equation˜1. The output of each PGNN block is computed as the sum of the shaped linear transformation and the correction path. In recursive simulations, PGNN blocks were applied iteratively for 20 steps to assess convergence behavior. For generalization gap analysis, we computed the mean squared error (MSE) between training and validation activations at each hidden layer. Subspace overlap was quantified using cosine squared principal angles. The contraction metric was defined as the mean absolute change between consecutive recursive outputs.
All models were trained using the Adam optimizer with a learning rate of $10^{-3}$ and default PyTorch initialization. Evaluation metrics included training loss, test accuracy, gradient norm tracking, and robustness under input perturbations. All figures in Sections VI and VII were generated directly from these setups.
VII-B Sensitivity to Input Alignment
@@
Observation In the aligned setting, structured networks converge significantly faster and exhibit lower gradient variance. In the randomized setting, convergence is slower for both, but the structured network still stabilizes earlier.
@@
VII-C Projection Incompleteness and Correction Load
@@
VII-D Stability Under Gradient Perturbation
@@
VII-E Pathway Decoupling Analysis
@@
VIII Task-Level Comparison on Fashion-MNIST
To examine how the proposed structured-corrective architecture behaves in practice, we compare it against a MLP on a real-world image classification task. This section does not aim to achieve state-of-the-art accuracy, but instead focuses on emergent training behavior and reliability across runs.
VIII-A Setup
We consider the Fashion-MNIST (FMNIST) dataset with two-layer networks using either (i) standard MLP layers or (ii) PGNN blocks as introduced earlier. Both models are trained for 20 epochs using the Adam optimizer and identical hyperparameters. The PGNN model employs an identity-based shaping operator and a learned correction path, while the MLP baseline is purely learned.
VIII-B Non-identity S on FMNIST
Beyond the identity-structured variant, we evaluate a fixed low–pass shaping operator inside the PGNN blocks:
$S\;=\;C^{\top}\operatorname{diag}(L)\,C,\qquad C\in\mathbb{R}^{d\times d}\ \text{is orthonormal DCT-II},\quad L_{i}=\mathbf{1}[\,i\leq 0.25d\,].$
All other settings (width, optimizer, batch size, seeds) are identical to the S=IS{=}I runs.
Figures˜4, 5 and 6 update our FMNIST plots to include this non-identity $S$. Across five seeds, PGNN with low-pass DCT (DCT-LP) closely tracks the identity variant: training loss and test accuracy curves are nearly overlapping, with a small early-epoch advantage and slightly tighter seed variance for DCT-LP; final accuracies differ by only a few tenths of a percent from the MLP baseline.
Figure 4: FMNIST test accuracy (mean $\pm$ sd over 5 seeds). Models: MLP, PGNN with identity S=IS{=}I, and PGNN with DCT low-pass $S$ (25% kept).
Figure 5: FMNIST training cross-entropy loss (mean over 5 seeds). The DCT-LP $S$ follows the identity case closely with a slight early-epoch advantage.
Figure 6: Initialization sensitivity (mean $\pm$ sd over 5 seeds). Shaded bands show seed variance. The non-identity $S$ yields similar or marginally lower variance than the identity case.
VIII-C Mechanistic diagnostics on FMNIST
[…]
VIII-D Interpretation
@@
Protocol. Gradient norms are total $\ell_{2}$ across parameters per epoch; the Jacobian spectrum is computed w.r.t. inputs at the logits on a single held-out batch and averaged over 32 samples.
[…]
PGNN and the MLP exhibit the same monotonic decline in total gradient norm, but PGNN stabilizes at a slightly lower level after the first few epochs (Fig. 7, left). The Jacobian spectrum on a held-out batch shows a characteristic pattern: PGNN has larger leading singular values yet a faster decay across the tail (right). This indicates stronger sensitivity along a few coherent directions with suppressed high-order/anisotropic responses—consistent with the structured path acting as a spectral shaper.
[…]
We evaluate centered linear CKA and Subspace-Overlap SOVkk on FMNIST using real activations (5 seeds, one held-out batch; k=16k{=}16 top PCs). CKA increases from L1 to L2 for all models, with a small gain for PGNN (S=I) at L2 (+0.005 over MLP), while SOV16{16} remains flat around 0.13 across layers and models; PGNN (DCT-LP) behaves comparably.
[…]
Model L1 L2
MLP 0.576 $\pm$ 0.008 0.641 $\pm$ 0.010
PGNN (S=I) 0.562 $\pm$ 0.008 0.646 $\pm$ 0.021
PGNN (DCT-LP) 0.569 $\pm$ 0.016 0.636 $\pm$ 0.027
[…]
Model L1 L2
MLP 0.130 $\pm$ 0.012 0.131 $\pm$ 0.005
PGNN (S=I) 0.131 $\pm$ 0.022 0.125 $\pm$ 0.012
PGNN (DCT-LP) 0.128 $\pm$ 0.010 0.132 $\pm$ 0.010
Table III: FMNIST layer-wise alignment (5 seeds). Same training protocol as §VIII.
IX Conclusion